OpenStudy (anonymous):

A 25 newton force applied on an object moves it 50 meters. The angle between the force and displacement is 40.0°. What is the value of work being done on the object? A. 0 joules B. 9.6 × 102 joules C. 1.2 × 103 joules D. -9.6 × 102 joules E. -1.2 × 103 joules

OpenStudy (anonymous):

\[W = F \Delta d \cos \theta\]where F = 25N \Delta d = 50m \theta = 40.0

OpenStudy (anonymous):

thank you!@ospreytriple

OpenStudy (anonymous):

You're welcome